OpenStudy (anonymous):

Find the solution set: 5x - 4 ≥ 26 or 29 - 3x > 2 a. {x|6≤x<9} b. {x|x≤6 or x>9} c. {real numbers} d. {x|x≠9}

OpenStudy (anonymous):

Set up each as and solve for x. I will do the first equality \[ \huge 5x - 4 \geq 26 \] \[ \huge 5x \geq 26 + 4 \] \[ \huge \frac{5x}{5} \geq \frac{30}{5} \] \[ \huge x \geq \frac{30}{5} \] \[ \huge x \geq 6 \] Now we know x is great or equal to 6. And we have two answer that stat that. You need to do the other one to find the answer. Do the next one equality \[ \huge 29 - 3x > 2 \]

OpenStudy (anonymous):

thank you!!!

OpenStudy (tabdo1226):

hi good afternoon

OpenStudy (anonymous):

hi!! :D